OpenStudy (vcabral1134):

I need help will give medal and fan!! 5x^3-3 if x<-7 f(x)= 2x^2-3 if -7<_ x<8 evaluate f(x) for x=-7 6+2x if x>_8

OpenStudy (anonymous):

@AlexanderPhantomhive @Firez @hartnn

OpenStudy (vcabral1134):

@KyanTheDoodle @Michele_Laino @texaschic101

OpenStudy (vcabral1134):

I need help can someone explain this to me?

OpenStudy (michele_laino):

there is a jump type discontinuity at x=7

OpenStudy (vcabral1134):

ok what else? I don't understand, I have to find a number @Michele_Laino

OpenStudy (michele_laino):

oops.. I meant at x=-7

OpenStudy (michele_laino):

we can compute the jump of the function at x = -7, as below: \[\begin{gathered} h\left( { - 7} \right) = \mathop {\lim }\limits_{x \to - 7 + } f\left( x \right) - \mathop {\lim }\limits_{x \to - 7 - } f\left( x \right) = \hfill \\ \hfill \\ = 2 \cdot {\left( { - 7} \right)^2} - 3 - 5 \cdot {\left( { - 7} \right)^3} + 3 \hfill \\ \end{gathered} \]

OpenStudy (vcabral1134):

ok

OpenStudy (vcabral1134):

do i solve the equation that you provided

OpenStudy (vcabral1134):

@TheSmartOne

OpenStudy (vcabral1134):

is the answer 1,807 ?? @Michele_Laino

OpenStudy (michele_laino):

I got 1,813

OpenStudy (vcabral1134):

ok maybe I did something wrong

OpenStudy (michele_laino):

we can write this: \[\begin{gathered} h\left( x \right) = \left( {2{x^2} - 3} \right) - \left( {5{x^3} - 3} \right) = \hfill \\ \hfill \\ = 2{x^2} - 3 - 5{x^3} + 3 = 2{x^2} - 5{x^3} \hfill \\ \end{gathered} \] now, please try to substitute \(x=-7\), what do you get?

OpenStudy (vcabral1134):

2* (49) - 3 - 5 * (-343) + 3 98 - 3 - (-1715) +3 95 - (-1712) 1807

OpenStudy (vcabral1134):

@Michele_Laino

OpenStudy (michele_laino):

we have: 98-3+1715+3=...?

OpenStudy (vcabral1134):

1813

OpenStudy (vcabral1134):

1813 @Michele_Laino

OpenStudy (michele_laino):

correct!

OpenStudy (vcabral1134):

Ok thank you so much!